solve each system by addition. State whether it's inconsistent or dependent.
2x + 3y =1
5x + 3y = 16

x + 5y = 10
-2x - 10y = 20

Answer by stanbon(75887)

solve each system by addition. State whether it's inconsistent or dependent.
1st: 2x + 3y =1
2nd: 5x + 3y = 16
Subtract 1st from 2nd to get:
3x=15
x=5
Substitute into 1st to solve for y, as follows:
2*5+3y=1
3y=-9
y=-3
Solution: x=5,y=-3; consistent
------------------------
1st: x + 5y = 10
2nd: -2x - 10y = 20
Multiply 1st by 2 to get:
3rd 2x+10y=20
Add 2nd and 3rd to get:
0=40
This says the x's in both equations and the y's in both equations
do not have the same meaning. The system is inconsistent.
---------------
Cheers,
Stan H.